# A Reinforcement Learning based Motion Planner for Quadrotor Autonomous Flight in Dense Environment

### v2

@@ V-C Sim to real transfer
We first train the policy in a customized Gym-based environment before transferring it to Gazebo simulation with PX4 firmware and then testing it on our real quadrotor platform (IMP250) with a Pixhawk 6c flight controller, without fine-tuning. We meticulously identified the Iris drone and our quadrotor models, achieving a relatively accurate representation. To ensure compatibility with low-level control, we utilize the PX4 multicopter rate controller and mixer to compute thrust and torques from the desired body rate and collective thrust outputs of the RL policy.
Additionally, we apply domain randomization during RL training to enhance policy robustness and prevent overfitting, compensating for numerical integration errors and modeling discrepancies between simulation and reality. We randomized the quadrotor’s physical properties, including mass, inertia, and thrust mapping, and ignored battery voltage drop and motor heating effects due to short flight times.
VI Results
@@ VI-A Training details
| [-] | 10 | [-] | -350 | |
| [-] | 0.01 | [-] | 1.5 | |
| [m] | 2.5 | [m] | 0.5 | |
| [s] | 0.02 | [m] | 0.3 | |
| [m] | 0.15 | [-] | 0.99 | |
| batch size [-] | 64 | learning rate [-] | 0.0003 |
The quadrotor’s physical properties and the RL algorithm’s hyperparameters are listed in Table I. We utilize the Iris quadrotor model from the PX4 firmware, adjusting its mass to 0.535 kg and the torque constant to 0.12 for our purposes. Following previous research, we employ Runge-Kutta 4th order numerical integration to update the quadrotor’s state, excluding aerodynamic drag during training due to its insignificance for the current .
@@ VI-B Analysis and Comparison
We compare our method with Fast-Planner [6] in terms of time to reach the same waypoint, energy consumption (integral of jerk), and computation time, with plans to include additional baselines in future work. Fast-Planner and our method employ ESDF for path searching and motion or trajectory generation. To ensure fairness, we remove control disturbances and initialize both methods from the same state. Our evaluation is conducted in two densely obstructed indoor scenarios, using identical models to assess and compare performance against Fast-Planner. The maximum velocity and acceleration for baseline methods are set to be and , respectively. For collision detection in the configuration space, the baseline methods use the same inflated obstacle size as ours, . Additionally, we impose stricter limits on z-axis position (), sensing horizon (), and searching horizon () to align with our method. Other parameters are set to default values.
| Time Span () | Energy () | |
| (1)Fast-Planner | 8.26 | 176.86 |
| (1)Ours | 5.80 | 935.88 |
| (2)Fast-Planner | 6.03 | 44.36 |
| (2)Ours | 4.78 | 576.76 |
To date, our method has demonstrated better performance in terms of time span, although the energy consumption remains much higher than that of Fast-Planner due to the aggressive control commands generated by the RL policy. Additionally, our method is less risk-aware compared to Fast-Planner, as it does not incorporate active perception, highlighting a potential direction for future work.
@@ VII Conclusion
In this letter, we propose a novel quadrotor motion planning framework that integrates path searching and motion generation using visibility-based methods and reinforcement learning. Our approach employs visibility path searching to generate safe waypoints with minimal path length, which are further optimized by an RL policy to produce low-level motion commands. We validate our method in two simulated indoor environments and compare it with Fast-Planner, demonstrating better performance in terms of time span.
Future work will focus on optimizing the RL policy to reduce energy consumption and enhance risk awareness through active yaw planning. We also plan to extend our method to real-world scenarios and evaluate its performance on a real quadrotor platform soon. Additionally, we aim to explore the potential of our approach in dynamic environments.
Acknowledgment
The authors would like to Yunlong Song for suggestions and discussions on RL policy training and sim to real transfer, Jia Xu and Zeshuai Chen for discussion on control and real-world platform.

### v4

@@ V-C Sim-to-real transfer
We first train the policy in a customized Gym-based environment before transferring it to Gazebo simulation with PX4 firmware and then testing it on our quadrotor platform without fine-tuning. We meticulously identified the Iris drone and our quadrotor models, achieving a relatively accurate representation. To ensure compatibility with low-level control, we utilize the PX4 multicopter rate controller and control allocator to compute thrusts from the desired body rate and collective thrust outputs of the RL policy.
Additionally, we apply domain randomization and action delay queue during RL training to enhance policy robustness and prevent overfitting, compensating for numerical integration errors and modeling discrepancies between simulation and reality. We randomized the quadrotor physical properties, including mass, inertia, and thrust mapping, and ignored battery voltage drop and motor heating effects due to short flight duration.
VI Experiment Results
| Time Span () | Energy () | |||||||
| Method | scene1 | scene2 | scene3 | scene4 | scene1 | scene2 | scene3 | scene4 |
| Fast-Planner[10] | 7.43 | 6.03 | 5.33 | 176.55 | 44.36 | 0 | 151.81 | |
| EGO-Planner[15] | 4.64 | 4.88 | 4.24 | 824.03 | 497.84 | 0 | 578.94 | |
| RESC | 4.5 | 4.48 | 11.7 | 4.00 | 697.11 | 682.53 | 1147.1 | 611.52 |
@@ VI-A Training details
The quadrotor physical properties and the RL algorithm’s hyperparameters are listed in Table II. Following previous research, we employ Runge-Kutta 4th order integration to update the quadrotor state, excluding aerodynamic drag during training due to its insignificance for the current .
@@
| [-] | 5.0 | [-] | 0.01 | |
| [-] | -600 | [-] | 300 | |
| [m] | 0.5 | [m] | 0.2 | |
| [s] | 0.02 | [-] | -1.5 | |
| [-] | 2048 | [-] | 0.99 | |
| batch size [-] | 128 | learning rate [-] | 0.0003 | |
| 10 | (GAE) | 0.95 |
@@ VI-B Analysis and Comparison
We compare our method with Fast-Planner [10] and EGO-Planner [15] in terms of time to reach the same position in four different scenes(Fig. 6). To ensure fairness, we remove control disturbances and initialize both methods from the same state. The baseline methods use the same maximum velocity (), acceleration (), and inflated obstacle size () as our method. Additionally, we impose stricter limits on z-axis position (), sensing horizon (), and searching horizon () to align with our method. Other parameters are set to default values. The comparison of baseline methods and our method is shown in Table I. In scenes 1, 2, and 4, our method outperforms the baseline approaches in terms of time efficiency, demonstrating its ability to generate more aggressive control commands. The drone achieves the desired motion accurately after executing the control commands, benefiting from the integration of dynamics in our policy network. In scene 3, only our method successfully reaches the goal, while the baseline methods fail to achieve the same result under the given acceleration and velocity constraints.
Consequently, our method shows better performance in terms of time efficiency, although it results in higher energy consumption compared to the baseline methods, primarily due to the aggressive control commands generated by the RL policy.
VI-C Simulation and Real-world Experiments
We present several experiments conducted in both simulation and real-world environments to evaluation the feasibility and performance of the proposed method. In simulation, we focus on navigating a quadrotor through some cluttered and randomized environments, where the RL policy generates control inputs that drive the quadrotor from the start point to the goal with control points provided by proposed searching method while avoiding obstacles. In the simulation environment, as shown in Fig. 8, the quadrotor follows control commands generated by RL policy, demonstrating the feasibility and real-time performance of our method. The actual trajectory driven by the RL policy is visualized in red.
To validate the performance of proposed planning system in real-world conditions, we conducted experiments in an indoor environment. The quadrotor’s pose data is provided by the NOKOV motion capture system, and the ESDF map is generated using the pose and depth images. As depicted in Fig. 9, the quadrotor successfully navigates around obstacles and adjusts its attitude to reach the goal. In Fig. 10, we show the actual trajectory of the quadrotor during a flight, further confirming the effectiveness of the RL policy in guiding the quadrotor to the goal while avoiding obstacles.
These experiments demonstrate the robustness and adaptability of the RL policy across both simulated and real-world environments, highlighting its potential for practical deployment in autonomous quadrotor motion planning tasks.
@@ VII Conclusion
In this letter, we present an improved framework for quadrotor local planning that integrates path searching and control generation using visibility-based methods and reinforcement learning. Our approach utilizes visibility path searching to generate safe control points with minimal path length, which are subsequently optimized by an RL policy to produce desired rate and throttle commands. We validate our method in several simulated environments and compare its performance against baseline approaches, showing superior time efficiency. We further conduct real-world experiments to demonstrate the feasibility and effectiveness of our method in practical applications.
Future work will focus on optimizing the RL policy to reduce energy consumption, enhance risk awareness through active yaw planning, and improve the success rate. Additionally, we aim to extend our method to outdoor real-world scenarios and expand it to fully 3D space. We also plan to investigate the potential of our approach in highly dynamic environments.